What is the angle between the hour and minute hand when it is 4:20?
Answer by lwsshak3(11628)

What is the angle between the hour and minute hand when it is 4:20?
**
The minute hand makes one revolution or 360º every hour.
At 4:00, the minute hand is at 12, its starting point=0º at 12
At 4:20, the minute hand traveled 20/60 or 1/3 of an hour which equates to 1/3 of 360º=120º from the starting point.
..
The hour hand makes one revolution or 360º every 12hours or 30º every hour
At 4:00, the hour hand is at 4, its starting point=120º from 12
At 4:20, the hour hand also traveled 1/3 of an hour which equates to 1/3 of 30º=10º
..
At 4:20, the minute hand ended at 120º from 12 and the hour hand 120+10=130º from 12.
So the angle between the hour and minute hand when it is 4:20=130º-120º=10º